OpenStudy (anonymous):

solve using the elimination method -7x -6y = -42; 5x - 2y = -14

OpenStudy (anonymous):

Make something the same, I chose the x terms. \[(5)*(-7x-6y=-42)\]\[(7)*(5x-2y=-14)\] Add \[-35x-30y=-210\]\[35x-14y=-98\] Wallah \[-44y=-308\]\[y=7\]

OpenStudy (anonymous):

Ok thanks, I got y =7 but I have to find out what x equals. but I get -7x equals 0. so what does x equal?

OpenStudy (anonymous):

Well, what do you have to multiply -7 by to get 0?

OpenStudy (anonymous):

0? isn't it?

OpenStudy (anonymous):

Ya.

OpenStudy (anonymous):

ok thanks